# Arithmetic of cuts in ordered abelian groups and of ideals over valuation rings

### v2

@@ 2.6. Ordered abelian groups modulo convex subgroups
For each subset $M$ of $G$ the preimage of $M/H$ under $\varphi_{H}$ is $M+H$. If $M$ is a final segment of $G$, then so is $M+H$.
@@
4): Since ${\mathcal{G}}(S)\subsetneq H$, we have $S+H\neq S$, so there is some $\alpha\in S+H$ such that $\alpha<S$ (since $S$ is a final segment). As $\alpha\in S+H$, it follows that $\alpha/H\in S/H$. We wish to show that $\alpha/H$ is the smallest element of $S/H$. Take any $\beta\in G$ such that $\beta/H<\alpha/H$. This means that $\beta+H<\alpha<S$, so $S\cap(\beta+H)=\emptyset$ and therefore, $\beta/H\notin S/H$. This proves our claim.
5): Take $\gamma\in G$ such that $\gamma/{\mathcal{G}}(S)$ is the smallest positive element of $G/{\mathcal{G}}(S)$ and therefore, $(\gamma/{\mathcal{G}}(S))\mathbb{Z}$ is a convex subgroup of $G/{\mathcal{G}}(S)$ and we have $(\gamma/{\mathcal{G}}(S))^{-}=(0/{\mathcal{G}}(S))^{+}$. Suppose that $S/{\mathcal{G}}(S)$ is nonprincipal. Then $(\gamma/{\mathcal{G}}(S))^{-}+S/{\mathcal{G}}(S)=S/{\mathcal{G}}(S)$, which implies that $(\gamma/{\mathcal{G}}(S))\mathbb{Z}\subseteq{\mathcal{G}}(S/{\mathcal{G}}(S)$, contradicting part 2) of our proposition. ∎
@@
We have that ${\mathcal{G}}(S)^{+}/{\mathcal{G}}(S)=(0/{\mathcal{G}}(S))^{+}$. Hence it follows from (10) together with part 3) of Lemma 2.6 that $S/{\mathcal{G}}(S)+{\mathcal{G}}(S)^{+}/{\mathcal{G}}(S)\neq S/{\mathcal{G}}(S)$ if and only if $S/{\mathcal{G}}(S)$ is principal. If inequality holds, then we also must have $S+{\mathcal{G}}(S)^{+}\neq S$. Now assume that equality holds. Then it must also hold for the respective preimages of $S/{\mathcal{G}}(S)$ and ${\mathcal{G}}(S)^{+}/{\mathcal{G}}(S)$ under the epimorphism $G\rightarrow G/{\mathcal{G}}(S)$. The preimage of $S/{\mathcal{G}}(S)$ is $S+{\mathcal{G}}(S)=S$. By part 5) of Lemma 2.14 with $\alpha=0$, ${\mathcal{G}}({\mathcal{G}}(S)^{+})={\mathcal{G}}(S)$, which shows that the preimage of ${\mathcal{G}}(S)^{+}/{\mathcal{G}}(S)$ is ${\mathcal{G}}(S)^{+}+{\mathcal{G}}(S)={\mathcal{G}}(S)^{+}$. Therefore, the equality $S/{\mathcal{G}}(S)+{\mathcal{G}}(S)^{+}/{\mathcal{G}}(S)=S/{\mathcal{G}}(S)$ implies the equality $S+{\mathcal{G}}(S)^{+}=S$. ∎
@@
Remark 2.28.
If $0\in\Lambda^{L}$, then ${\mathcal{G}}(\Lambda)\subseteq\Lambda^{L}$ by part 4) of Lemma 2.13. Symmetrically, if $0\in\Lambda^{R}$, then ${\mathcal{G}}(\Lambda)\subseteq\Lambda^{R}$. It follows that ${\mathcal{G}}(\Lambda)=G$ if and only if $\Lambda^{L}$ or $\Lambda^{R}$ is empty, that is, if $\Lambda$ is not a Dedekind cut.
Theorem 2.29.
@@
The quasi-cut $\Lambda/H$ is a cut in $G/H$ if and only if $\Lambda^{L}/H\cap\Lambda^{R}/H=\emptyset$. This holds if and only if $\Lambda^{L}+H\subseteq\Lambda^{L}$. By definition of ${\mathcal{G}}(\Lambda)$, this in turn holds if and only if $H\subseteq{\mathcal{G}}(\Lambda)$.
The last statement follows from the definition of ${\mathcal{G}}(\Lambda)$ together with Proposition 2.26. ∎
@@ 2.7. The equation $S_{1}+T=S_{2}$ in ordered abelian groups of arbitrary rank
Take final segments $S_{1},S_{2},T_{0}$ and a proper convex subgroup $H$ of $G$. If $\bar{T}=T_{0}/H$ is the largest solution of $S_{1}/H+\bar{T}=S_{2}/H$, then its preimage $T=T_{0}+H$ is the largest solution of $S_{1}+T=S_{2}\,$.
@@
to be the preimage of $\widehat{T/{\mathcal{G}}(T)}$ under $\varphi_{{\mathcal{G}}(T)}:G\rightarrow G/{\mathcal{G}}(T)$. Note that $T^{\diamondsuit}$ is a final segment of $G$. Observe that $T^{\diamondsuit}=T$ holds if and only if $T/{\mathcal{G}}(T)$ is closed in $G/{\mathcal{G}}(T)$. If ${\mathcal{G}}(T)$ is trivial, then $T^{\diamondsuit}=\widehat{T}$. We view $T^{\diamondsuit}$ as the “deep closure” of $T$.
@@
2) For each $\alpha\in G$, the preimage of $(\alpha/{\mathcal{G}}(T))^{-}$ under $\varphi_{cG(T)}$ is $a^{-}+{\mathcal{G}}(T)$.
@@
1): By part 1) of Lemma 2.17, ${\mathcal{G}}(-T^{c})={\mathcal{G}}(T^{c})={\mathcal{G}}(T)$, hence the preimage of $-T^{c}/{\mathcal{G}}(T)$ under $\varphi_{{\mathcal{G}}(T)}$ is $(-T^{c})+{\mathcal{G}}(T)=-T^{c}$. It remains to show that $-T^{c}/{\mathcal{G}}(T)=-(T/{\mathcal{G}}(T))^{c}$. Since ${\mathcal{G}}(T^{c})={\mathcal{G}}(T)\,$, the sets $S(T)/{\mathcal{G}}(T)$ and $T^{c}/{\mathcal{G}}(T)$ are disjoint. On the other hand, their union is $G/{\mathcal{G}}(T)\,$, so $T^{c}/{\mathcal{G}}(T)=(T/{\mathcal{G}}(T))^{c}$, showing that $-T^{c}/{\mathcal{G}}(T)=-(T/{\mathcal{G}}(T))^{c}$ holds.
@@
3): If there is $\alpha\in G$ such that $\alpha/{\mathcal{G}}(T)$ is the infimum of $T/{\mathcal{G}}(T)$ in $G/{\mathcal{G}}(T)$, then $\widehat{T/{\mathcal{G}}(T)}=(\alpha/{\mathcal{G}}(T))^{-}$, whose preimage in $G$ is $\alpha^{-}+{\mathcal{G}}(T)$. Otherwise, $\widehat{T/{\mathcal{G}}(T)}=T/{\mathcal{G}}(T)$, whose preimage is $T+{\mathcal{G}}(T)=T$. ∎
@@
2): We observe that ${\mathcal{G}}(S_{2}/{\mathcal{G}}(S_{2}))$ is trivial by part 2) of Proposition 2.26. By assumption, $S_{1}/{\mathcal{G}}(S_{2})$ is nonprincipal, so we can apply Lemma 2.22 to infer that $\overline{T}=\widehat{T_{1}/{\mathcal{G}}(S_{2})}$ is the largest final segment of $G/{\mathcal{G}}(S_{2})$ such that $S_{1}/{\mathcal{G}}(S_{2})+\overline{T}=S_{2}/{\mathcal{G}}(S_{2})$. Since $S_{1}+T_{0}=S_{2}$, we know from part 2) of Lemma 2.14 that ${\mathcal{G}}(S_{2})={\mathcal{G}}(S_{1}+T_{0})={\mathcal{G}}(S_{1})\cup{\mathcal{G}}(T_{0})$, whence ${\mathcal{G}}(T_{0})\subseteq{\mathcal{G}}(S_{2})$. Using part 4) of the same lemma, we conclude that ${\mathcal{G}}(T_{1})={\mathcal{G}}(T_{0}+{\mathcal{G}}(S_{2}))={\mathcal{G}}(T_{0})\cup{\mathcal{G}}(S_{2})={\mathcal{G}}(S_{2})$. Therefore, from its definition and Lemma 2.30 we obtain that $T_{1}^{\diamondsuit}$ is the largest final segment $T$ of $G$ such that $T/{\mathcal{G}}(S_{2})=\widehat{T_{1}/{\mathcal{G}}(S_{2})}$. This proves statement 2). ∎
@@
Since ${\mathcal{G}}(S+S)={\mathcal{G}}(S)$ by Corollary 2.15 and $S+{\mathcal{G}}(S)=S$, we can apply the foregoing proposition to $T_{1}=S$. If $S/{\mathcal{G}}(S)$ is nonprincipal, then our claim follows from part 2) of Proposition 2.32. Now assume that $S/{\mathcal{G}}(S)$ is principal. Then by part 1) of Proposition 2.32, $S$ is the largest solution, and we need to show that $S=S^{\diamondsuit}$ in this case. As $S/{\mathcal{G}}(S)$ is principal, it is closed in $G/{\mathcal{G}}(S)$ and we indeed have $S^{\diamondsuit}=S$ by our observation before Lemma 2.31. ∎
@@
To simplify notation, we set ${\mathcal{G}}_{2}:={\mathcal{G}}(S_{2})$. Assume that condition (*) holds. From part 2) of Proposition 2.26 we know that ${\mathcal{G}}(S_{1}/{\mathcal{G}}_{2})={\mathcal{G}}(S_{2}/{\mathcal{G}}_{2})=\{0\}$. Hence by Proposition 2.24, the largest solution of $S_{1}/{\mathcal{G}}_{1}+\bar{T}=S_{2}/{\mathcal{G}}_{1}$ is $\bar{T}=S_{2}/{\mathcal{G}}_{2}-\alpha/{\mathcal{G}}_{2}=(S_{2}-\alpha)/{\mathcal{G}}_{2}$ if $S_{1}/{\mathcal{G}}_{2}=(\alpha/{\mathcal{G}}_{2})^{-}$, and $\bar{T}=(S_{2}/{\mathcal{G}}_{2}-(S_{1}/{\mathcal{G}}_{2})^{c})\,\widehat{}$ if $S_{1}/{\mathcal{G}}_{2}$ is nonprincipal.
By Lemma 2.30 the preimage $T$ of $\bar{T}$ under the epimorphism $G\rightarrow G/{\mathcal{G}}_{2}$ is the largest solution of (1). As stated in the introduction, $S_{1}-_{\rm fs}S_{2}$ is the largest solution of $S_{1}+T\subseteq S_{2}$, so it follows that $T=S_{1}-_{\rm fs}S_{2}\,$.
If $\bar{T}=(S_{2}-\alpha)/{\mathcal{G}}_{2}\,$, then $T=(S_{2}-\alpha)+{\mathcal{G}}_{2}=S_{2}-\alpha$ by part 3) of Lemma 2.14. Now assume that $S_{1}/{\mathcal{G}}_{2}$ is nonprincipal and $\bar{T}=(S_{2}/{\mathcal{G}}_{2}-(S_{1}/{\mathcal{G}}_{2})^{c})\,\widehat{}$. By part 4) of Proposition 2.26, the former implies that ${\mathcal{G}}_{2}\subseteq{\mathcal{G}}(S_{1})$, whence ${\mathcal{G}}(S_{1})={\mathcal{G}}_{2}\,$. Therefore, the preimage of $-(S_{1}/{\mathcal{G}}_{2})^{c}$ is $-S_{1}^{c}$ by part 1) of Lemma 2.31. By Lemma 2.25, the preimage of $S_{2}/{\mathcal{G}}_{2}$ is $S_{2}+{\mathcal{G}}_{2}=S_{2}\,$, hence the preimage of $S_{2}/{\mathcal{G}}_{2}-(S_{1}/{\mathcal{G}}_{2})^{c}$ is $S_{2}-S_{1}^{c}$. Hence the preimage of $\bar{T}$ is $T=(S_{2}-S_{1}^{c})^{\diamondsuit}$.
Now assume that condition (*) does not hold. Then one of the cases i) or ii) must hold; observe that if ${\mathcal{G}}(S_{1})\subsetneq{\mathcal{G}}_{2}$, then by part 4) of Proposition 2.26, $S_{1}/{\mathcal{G}}_{2}$ is principal and condition (*) holds. Note that in both cases, $S_{1}$ is nonprincipal. In case i), Proposition 2.16 shows that equation (1) has no solution. Assume that case ii) holds and that the final segment $T$ is a solution of equation (1). Then $S_{1}/{\mathcal{G}}_{2}+T/{\mathcal{G}}_{2}=S_{2}/{\mathcal{G}}_{2}$, but this contradicts part 2) of Lemma 2.4 applied to the group $G/{\mathcal{G}}_{2}\,$. ∎
@@
1) Assume that ${\mathcal{G}}(S_{2})\subsetneq{\mathcal{G}}(S_{1})$. Then there exists an element $\alpha\in S_{2}$ such that $S^{\prime}_{2}:=(\alpha+{\mathcal{G}}(S_{1}))^{+}\subsetneq S_{2}$ is the largest final segment contained in $S_{2}$ having invariance group ${\mathcal{G}}(S_{1})$. Further, $T=S_{2}^{\prime}-_{\rm fs}S_{1}$ is the largest solution of $S_{1}+T=S_{2}^{\prime}$, and it is equal to $S_{2}-_{\rm fs}S_{1}$.
2) Assume that ${\mathcal{G}}(S_{1})={\mathcal{G}}(S_{2})$ and $S_{1}/{\mathcal{G}}(S_{1})$ is nonprincipal, but $S_{2}/{\mathcal{G}}(S_{1})$ is principal, say $S_{2}/{\mathcal{G}}(S_{1})=(\alpha/{\mathcal{G}}(S_{1}))^{-}$ for some $\alpha\in S_{2}\,$. Then $S^{\prime}_{2}:=(\alpha+{\mathcal{G}}(S_{1}))^{+}\subsetneq S_{2}$ is the largest final segment contained in $S_{2}$ that is nonprincipal. Again, $T=S_{2}^{\prime}-_{\rm fs}S_{1}$ is the largest solution of $S_{1}+T=S_{2}^{\prime}$, and it is equal to $S_{2}-_{\rm fs}S_{1}$.

### v3

@@ 2.6. Ordered abelian groups modulo convex subgroups
For each subset $M$ of $G$ the preimage $\varphi_{H}^{-1}(M/H)$ of $M/H$ under $\varphi_{H}$ is $M+H$. If $M$ is a final segment of $G$, then so is $M+H$.
@@
4): If ${\mathcal{G}}(S)\subsetneq H$, then we have $S+H\neq S$, so there is some $\alpha\in S+H$ such that $\alpha<S$ (since $S$ is a final segment). As $\alpha\in S+H$, it follows that $\alpha/H\in S/H$. We wish to show that $\alpha/H$ is the smallest element of $S/H$. Take any $\beta\in G$ such that $\beta/H<\alpha/H$. This means that $\beta+H<\alpha<S$, so $S\cap(\beta+H)=\emptyset$ and therefore, $\beta/H\notin S/H$. This proves our claim.
5): Take $\gamma\in G$ such that $\gamma/{\mathcal{G}}(S)$ is the smallest positive element of $G/{\mathcal{G}}(S)$ and therefore, $(\gamma/{\mathcal{G}}(S))\mathbb{Z}$ is a convex subgroup of $G/{\mathcal{G}}(S)$ and we have $(\gamma/{\mathcal{G}}(S))^{-}=(0/{\mathcal{G}}(S))^{+}$. Suppose that $S/{\mathcal{G}}(S)$ is nonprincipal. Then $(\gamma/{\mathcal{G}}(S))^{-}+S/{\mathcal{G}}(S)=(0/{\mathcal{G}}(S))^{+}+S/{\mathcal{G}}(S)=(0/{\mathcal{G}}(S))^{-}+S/{\mathcal{G}}(S)=S/{\mathcal{G}}(S)$ by equation (15), which implies that $(\gamma/{\mathcal{G}}(S))\mathbb{Z}\subseteq{\mathcal{G}}(S/{\mathcal{G}}(S))$, contradicting part 2) of our proposition. ∎
@@
We have that ${\mathcal{G}}(S)^{+}/{\mathcal{G}}(S)=(0/{\mathcal{G}}(S))^{+}$. Hence it follows from (10) together with part 3) of Lemma 2.6 that $S/{\mathcal{G}}(S)+{\mathcal{G}}(S)^{+}/{\mathcal{G}}(S)\neq S/{\mathcal{G}}(S)$ if and only if $S/{\mathcal{G}}(S)$ is principal. If inequality holds, then we also must have $S+{\mathcal{G}}(S)^{+}\neq S$. Now assume that equality holds. Then it must also hold for the respective preimages of $S/{\mathcal{G}}(S)$ and ${\mathcal{G}}(S)^{+}/{\mathcal{G}}(S)$ under $\varphi_{{\mathcal{G}}(S)}\,$. By Lemma 2.24, $\varphi_{{\mathcal{G}}(S)}^{-1}(S/{\mathcal{G}}(S))=S+{\mathcal{G}}(S)=S$. By part 5) of Lemma 2.13 with $\alpha=0$, ${\mathcal{G}}({\mathcal{G}}(S)^{+})={\mathcal{G}}(S)$, which shows that $\varphi_{{\mathcal{G}}(S)}^{-1}({\mathcal{G}}(S)^{+}/{\mathcal{G}}(S))={\mathcal{G}}(S)^{+}+{\mathcal{G}}(S)={\mathcal{G}}(S)^{+}$. Therefore, the equality $S/{\mathcal{G}}(S)+{\mathcal{G}}(S)^{+}/{\mathcal{G}}(S)=S/{\mathcal{G}}(S)$ implies the equality $S+{\mathcal{G}}(S)^{+}=S$. ∎
@@
Proposition 2.27.
@@
The quasi-cut $\Lambda/H$ is a cut in $G/H$ if and only if $\Lambda^{L}/H\cap\Lambda^{R}/H=\emptyset$. This holds if and only if $\Lambda^{R}+H\subseteq\Lambda^{R}$. By definition of ${\mathcal{G}}(\Lambda)$, this in turn holds if and only if $H\subseteq{\mathcal{G}}(\Lambda)$.
The last statement follows from the definition of ${\mathcal{G}}(\Lambda)$ together with part 2) of Proposition 2.25. ∎
@@ 2.7. The equation $S_{1}+T=S_{2}$ in ordered abelian groups of arbitrary rank
Take final segments $S_{1},S_{2},T_{0}$ and a proper convex subgroup $H$ of $G$. If $\overline{T}=\overline{T}_{0}$ is the largest solution of $S_{1}/H+\overline{T}=S_{2}/H$, then its preimage $T=\varphi_{H}^{-1}(\overline{T}_{0})$ is the largest solution of $S_{1}+T=S_{2}\,$. In particular, if $\overline{T}=T_{0}/H$ is the largest solution of $S_{1}/H+\overline{T}=S_{2}/H$, then its preimage $T=T_{0}+H$ under $\varphi_{H}$ is the largest solution of $S_{1}+T=S_{2}\,$.
@@
to be the preimage of $\widehat{T/{\mathcal{G}}(T)}$ under $\varphi_{{\mathcal{G}}(T)}\,$. Note that $T^{\diamondsuit}$ is a final segment of $G$. Observe that $T^{\diamondsuit}=T$ holds if and only if $T/{\mathcal{G}}(T)$ is closed in $G/{\mathcal{G}}(T)$. If ${\mathcal{G}}(T)$ is trivial, then $T^{\diamondsuit}=\widehat{T}$. It follows from part 6) of Lemma 2.12 that $T^{\diamondsuit}$ is closed also if ${\mathcal{G}}(T)$ is not trivial. We view $T^{\diamondsuit}$ as the “deep closure” of $T$.
@@
2) For each $\alpha\in G$, the preimage of $(\alpha/{\mathcal{G}}(T))^{-}$ under $\varphi_{{\mathcal{G}}(T)}$ is $a^{-}+{\mathcal{G}}(T)$.
@@
1): By part 1) of Lemma 2.16, ${\mathcal{G}}(-T^{c})={\mathcal{G}}(T^{c})={\mathcal{G}}(T)$, hence the preimage of $-T^{c}/{\mathcal{G}}(T)$ under $\varphi_{{\mathcal{G}}(T)}$ is $-T^{c}+{\mathcal{G}}(T)=-T^{c}$. It remains to show that $-T^{c}/{\mathcal{G}}(T)=-(T/{\mathcal{G}}(T))^{c}$. Since ${\mathcal{G}}(T^{c})={\mathcal{G}}(T)\,$, the sets $T/{\mathcal{G}}(T)$ and $T^{c}/{\mathcal{G}}(T)$ are disjoint. On the other hand, their union is $G/{\mathcal{G}}(T)\,$, so $T^{c}/{\mathcal{G}}(T)=(T/{\mathcal{G}}(T))^{c}$, showing that $-T^{c}/{\mathcal{G}}(T)=-(T/{\mathcal{G}}(T))^{c}$ holds.
@@
3): If there is $\alpha\in G$ such that $\alpha/{\mathcal{G}}(T)$ is the infimum of $T/{\mathcal{G}}(T)$ in $G/{\mathcal{G}}(T)$, then $\widehat{T/{\mathcal{G}}(T)}=(\alpha/{\mathcal{G}}(T))^{-}$, whose preimage under $\varphi_{{\mathcal{G}}(T)}$ is $\alpha^{-}+{\mathcal{G}}(T)$. Otherwise, $\widehat{T/{\mathcal{G}}(T)}=T/{\mathcal{G}}(T)$, whose preimage is $T+{\mathcal{G}}(T)=T$. ∎
@@
2): We observe that ${\mathcal{G}}(S_{2}/{\mathcal{G}}(S_{2}))$ is trivial by part 2) of Proposition 2.25. By assumption, $S_{1}/{\mathcal{G}}(S_{2})$ is nonprincipal, so we can apply Lemma 2.21 to infer that $\overline{T}=\widehat{T_{1}/{\mathcal{G}}(S_{2})}$ is the largest solution of $S_{1}/{\mathcal{G}}(S_{2})+\overline{T}=S_{2}/{\mathcal{G}}(S_{2})$. Therefore, from its definition and Lemma 2.28 we obtain that $T=T_{1}^{\diamondsuit}$ is the largest solution of (1). ∎
@@
We apply Proposition 2.30 with $T_{0}=S$. Since ${\mathcal{G}}(S+S)={\mathcal{G}}(S)$ by Corollary 2.14 and $S+{\mathcal{G}}(S)=S$, we obtain $T_{1}=S$. If $S/{\mathcal{G}}(S)$ is nonprincipal, then our claim follows from part 2) of Proposition 2.30. Now assume that $S/{\mathcal{G}}(S)$ is principal. Then by part 1) of Proposition 2.30, $S$ is the largest solution, and we need to show that $S=S^{\diamondsuit}$ in this case. As $S/{\mathcal{G}}(S)$ is principal, it is closed in $G/{\mathcal{G}}(S)$ and we indeed have $S^{\diamondsuit}=S$ by our observation before Lemma 2.29. ∎
@@
To simplify notation, we set ${\mathcal{G}}_{2}:={\mathcal{G}}(S_{2})$. Assume that condition ()(*)( ∗ ) holds. From part 2) of Proposition 2.25 we know that ${\mathcal{G}}(S_{1}/{\mathcal{G}}_{2})={\mathcal{G}}(S_{2}/{\mathcal{G}}_{2})=\{0\}$. Hence by Proposition 2.23, the largest solution of $S_{1}/{\mathcal{G}}_{1}+\overline{T}=S_{2}/{\mathcal{G}}_{1}$ is $\overline{T}=S_{2}/{\mathcal{G}}_{2}-\alpha/{\mathcal{G}}_{2}=(S_{2}-\alpha)/{\mathcal{G}}_{2}$ if $S_{1}/{\mathcal{G}}_{2}=(\alpha/{\mathcal{G}}_{2})^{-}$, and $\overline{T}=(S_{2}/{\mathcal{G}}_{2}-(S_{1}/{\mathcal{G}}_{2})^{c})\,\widehat{}$ if $S_{1}/{\mathcal{G}}_{2}$ is nonprincipal.
By Lemma 2.28 the preimage $T$ of $\overline{T}$ under $\varphi_{{\mathcal{G}}_{2}}$ is the largest solution of (1). As stated in the introduction, $S_{2}-_{\rm fs}S_{1}$ is the largest solution of $S_{1}+T\subseteq S_{2}$, so it follows that $T=S_{2}-_{\rm fs}S_{1}\,$.
If $\overline{T}=(S_{2}-\alpha)/{\mathcal{G}}_{2}\,$, then $T=(S_{2}-\alpha)+{\mathcal{G}}_{2}=S_{2}-\alpha$ by part 3) of Lemma 2.13. Now assume that $S_{1}/{\mathcal{G}}_{2}$ is nonprincipal and $\overline{T}=(S_{2}/{\mathcal{G}}_{2}-(S_{1}/{\mathcal{G}}_{2})^{c})\,\widehat{}$. By part 4) of Proposition 2.25, the former implies that ${\mathcal{G}}_{2}\subseteq{\mathcal{G}}(S_{1})$, whence ${\mathcal{G}}(S_{1})={\mathcal{G}}_{2}\,$. Therefore, $\varphi_{{\mathcal{G}}_{2}}^{-1}(-(S_{1}/{\mathcal{G}}_{2})^{c})=-S_{1}^{c}$ by part 1) of Lemma 2.29. By Lemma 2.24, $\varphi_{{\mathcal{G}}_{2}}^{-1}(S_{2}/{\mathcal{G}}_{2})=S_{2}+{\mathcal{G}}_{2}=S_{2}\,$, hence $\varphi_{{\mathcal{G}}_{2}}^{-1}(S_{2}/{\mathcal{G}}_{2}-(S_{1}/{\mathcal{G}}_{2})^{c})=S_{2}-S_{1}^{c}$. Thus the preimage of $\overline{T}$ is $T=(S_{2}-S_{1}^{c})^{\diamondsuit}$.
Now assume that condition ()(*)( ∗ ) does not hold. Then one of the cases i) or ii) must hold; observe that if ${\mathcal{G}}(S_{1})\subsetneq{\mathcal{G}}_{2}$, then by part 4) of Proposition 2.25, $S_{1}/{\mathcal{G}}_{2}$ is principal and condition ()(*)( ∗ ) holds. Note that in both cases, $S_{1}$ is nonprincipal. In case i), Proposition 2.15 shows that equation (1) has no solution. Assume that case ii) holds and that the final segment $T$ is a solution of equation (1). Then $S_{1}/{\mathcal{G}}_{2}+T/{\mathcal{G}}_{2}=S_{2}/{\mathcal{G}}_{2}$, but this contradicts part 2) of Lemma 2.4 applied to the group $G/{\mathcal{G}}_{2}\,$. ∎
@@
1) Assume that ${\mathcal{G}}(S_{2})\subsetneq{\mathcal{G}}(S_{1})$. Then there exists an element $\alpha\in S_{2}$ such that $S^{\prime}_{2}:=(\alpha+{\mathcal{G}}(S_{1}))^{+}\subsetneq S_{2}$ is the largest final segment contained in $S_{2}$ having invariance group ${\mathcal{G}}(S_{1})$.
2) Assume that ${\mathcal{G}}(S_{1})={\mathcal{G}}(S_{2})$ and $S_{1}/{\mathcal{G}}(S_{1})$ is nonprincipal, but $S_{2}/{\mathcal{G}}(S_{1})$ is principal, say $S_{2}/{\mathcal{G}}(S_{1})=(\alpha/{\mathcal{G}}(S_{1}))^{-}$ for some $\alpha\in S_{2}\,$. Then $S^{\prime}_{2}:=(\alpha+{\mathcal{G}}(S_{1}))^{+}\subsetneq S_{2}$ is the largest final segment contained in $S_{2}$ that is nonprincipal.
In both cases, condition ()(*)( ∗ ) of Theorem 2.33 is satisfied with $S^{\prime}_{2}$ in place of $S_{2}$ and $T=S_{2}^{\prime}-_{\rm fs}S_{1}$ is the largest solution of $S_{1}+T=S_{2}^{\prime}$. Further, $S_{2}^{\prime}-_{\rm fs}S_{1}=S_{2}-_{\rm fs}S_{1}$.